Crew 216 Astronomy Report 30Nov2019

[title Astronomy Report – November 30th]

[category astronomy-report]

Crew 216 Astronomy Report 30-NOV-19
Crew Astronomer: Michael Ho

MDRS ROBOTIC OBSERVATORY
Robotic Telescope Requested: none
Objects to be Imaged this Evening: none
Images submitted with this report: nil
Problems Encountered: nil

MUSK OBSERVATORY
Solar Features Observed: None.
Images submitted with this report: nil
Problems Encountered: Liner protecting edge of upper shutter is half-detached. Please see attached photo.

Crew 216 Science Report 30Nov2019

[title Science Report – November 30th]

[category science-report]

Crew 216 Research Report 30-Nov-2019
Sol 6
Submitted by: Whittle

1) Psych Scale (TAMU IRB2019-1042D) – Nothing further to report.

2) Medical Casts – Continuation of Crew 215 research. Crushed 60% of gypsum, will continue tomorrow and begin purifying.

3) In-Situ Fluorescent Mineral Prospecting – Planning reconnaissance EVA for 03-Dec-2019.

4) MDRS Mapping – Nothing further to report (awaiting return of map with desired annotations).

5) MDRS Communications – Completed GIS view shades of repeater sites to West and South of Hab. Attempted to climb ridge to North of Hab to find rebroadcast point, climb aborted due to steepness. Will complete GIS reconnaissance of same point and potentially try again later in mission.

Crew 216 GreenHab Report 30Nov2019

[title GreenHab Report – November 30th]

[category greenhab-report]

Crew 216 GreenHab Report 30-NOV-19
Crew GreenHab Officer: Michael Ho

Environmental control: Door opened for ventilation at 12.10 pm, closed at 5.15 pm.
Average temperature: 20C
Hours of supplemental light: Automatic system
Daily water usage for crops: 14L
Daily water usage for research and/or other purposes: nil
Water in Blue Tank – enough
Time(s) of watering for crops: 09.40 am; 4.10 pm.

SEASONAL PLANTING HISTORY:
(Established plants introduced since 03/11/2019 till 22/11/2019 please refer to Crew 215 GreenHab Report of 22/11/2019.)
Crew 216:
23/11/2019:
seeded Zucchini in 2 large black pots
seeded Chilli Pepper ‘Shishito’ in 2 2-gallon black pots
seeded Sweet Greenfield Pepper ‘California Wonder’ in 2 2-gallon black pots
seeded Swiss Chard ‘Flamingo Pink’ in 4 2-gallon black pots
transplanted three self-seeded tomato plants in large blue tub transplanted to 2-gallon pots
24/11/2019:
1 avocado seed in small black pot
26/11/2019:
1 avocado seed in small black pot

OTHER NOTES:
Extra wires for hanging trays 1xcurrent length; 2xshorter length
Harvest: None
Support/supplies requested: 5-10pcs plywood 13"x17"x3/16".

Commander Report – November 30th

Crew 216 Commander Report 30-NOV-2019
Sol: 6
Summary Title: Turning Point
Author: Marc Levesque
Crew 216 reached the halfway point for its mission with a sense of enjoyment and appreciation for the time spent so far at MDRS, but also a sense of sadness that there is only another week left. This morning provided an opportunity for some, especially me, to take a time out from the all the activity and reflect upon what the crew has accomplished so far and what still needs to be done in the remaining days. I knew coming into the mission that time would fly, and although some of the crew felt the first few days dragged, all now feel that the pace is accelerating towards our departure from the Red Planet. On today’s EVA we scouted for a route to a potential repeater site but were unable to reach the summit of North Ridge. In continuing our reconnaissance for possible routes, we traversed a landscape that made us truly feel like we were on Mars, something that made up for not reaching our objective. This evening I shared with the crew my very positive feelings about how well they have been doing, keeping the Hab clean and tidy, getting along with each other, and generally helping everyone with their projects. I feel deep gratitude for having been selected as commander of these talented, friendly, and bright individuals.

EVA Report – November 30th

Crew 216 EVA Report 30-Nov-2019
EVA #08
Author: Levesque

Purpose of EVA: Evaluate potential communications repeater site on North Ridge

Start time for EVA: 1330
End time for EVA: 1530

Narrative: The EVA team walked to the bottom of the northeast corner of the North Ridge to assess the viability of a safe route to the top that had been determined from cartographic analysis. Once again, however, the dictum was proven true: the map is not the territory. Upon reaching the northeast corner, the team determined that the steepness of the slope and the footing was not safe or suitable to pursue. In search of an alternative, the team continued west around the north side completing a loop without finding one. The team then returned to the Hab via a direct route. A future reconnaissance in drier conditions should be considered, as a GIS viewshed analysis of the North Ridge indicates it would be an ideal site for a radio repeater if a relatively easy route to the summit can be found.

EVA Participants: Levesque, Lawson, Kennedy, Alexandrova

Routes per MDRS Map:
1) Walk to the northeast corner of North Ridge at 518624E 4251655N
2) Continued west to 518248E 4251639N
3) Returned to Hab from this point

Mode of travel: Walk
Vehicles used: N/A

Operations Report – November 30th

Crew 216 Operations Report 30-Nov-19
SOL: 6
Name of person filing report: Mike Lawson

Non-nominal systems: RAM rover power charging stations inoperable, electrical issue, ongoing.
Notes on non-nominal systems:

Generator: run
Hours run: 11.5
From what time last night: 1902
To what time this morning: 0630
List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before the generator is run at night: 71%

Diesel Reading – 38
Station Propane Reading – 56
Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 47
Water Meter (units): 1463316
Water (static tank) (gal): 315
Static to Loft Pump used – yes
Water in Green Hab (gal): 263
Water in Science Dome (gal): 0

Toilet tank emptied: yes

Deimos rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used:
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: no
Hours: 114.5
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 100
Currently charging:

Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Curiosity rover used: no
Hours: 119.6
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 100
Currently charging: yes

Notes on rovers: will have to juggle the charging sequence.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:

HabCar used and why, where?
CrewCar used and why, where?

General notes and comments: EVA to North Ridge, circumnavigated hill mass. No obvious routes for the summit in sim. No recon of potential radio repeater site.

Summary of internet: No serious issues.
Summary of suits and radios: No issues.
Summary of Hab operations: Routine maintenance and housekeeping.
Summary of GreenHab operations: routine watering.
Summary of Science Dome operations: Gypsum processing continued.
Summary of RAM operations: RAM Rover power charging stations inoperable, electrical issues remain unresolved.
Summary of any observatory issues: Sun observations performed.

Journalist Report – November 30th

Crew 216 Journalist Report 30-NOV-19
Sol 6
Author: Evgenia Alexandrova

How does the time work on Mars?

Time flies when half of the mission is already behind.

Time is really long in the airlock during the five-minute depressurization.

Time flies when it is already the Capcom window and you haven’t written your report yet.

Time is really long when you wait for the bread to be cooked.

There is this film “Room” by Lenny Abrahamson in which a 9-years old boy, who was born and grew up in a 16 square meters room, finally gets out of this room and discovers the rest of the world. Excited, he shares with his mother, that the time in the rest of the world flies much faster than in the room because the same amount of time has to be spread on a much larger surface.

Following this logic, time on the EVA flies faster than in the HAB. So the two halves of the crew experience very different sensations during this moment.

All I know is that time is not linear, the greatest minds have proved it long ago.

Anonymous Haiku for this Saturday night:

Steady pulse of pump
Soft hum of battery bank
Voices talk of space

Sol Summary – November 30th

Crew 216 Sol Summary Report 30-NOV-19
Sol: 6
Summary Title: Rest and Reconnaissance
Author: Marc Levesque

Mission Status: Active

Sol Activity Summary: The crew spent the morning waiting for the ground to dry sufficiently to allow an afternoon EVA. This provided an opportunity for crew rest and pursuing projects for filming, the GreenHab, and astronomy. With sunny weather drying out the landscape, the afternoon EVA was approved and undertaken with caution. While the main objective of the EVA was to reach the summit of the North Ridge to determine a suitable repeater site was not achieved because of ground conditions, a reconnaissance and assessment of route alternatives to reach the top of the North Ridge was accomplished.

Look Ahead Plan: A film crew from Al Roker Entertainment will be arriving around 0900 to spend the day filming the MDRS, the crew, and its activities.

Anomalies in work: N/A.

Weather: Clearing and cold.

Crew Physical Status: Healthy

EVA: N/A

Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, EVA Report, Astronomy Report, and Photos.

Support Requested: None

Operations Report – November 29th

rew 216 Operations Report 29-Nov-19
SOL: 5
Name of person filing report: Mike Lawson
Non-nominal systems: RAM rover power charging stations inoperable, electrical issue, ongoing.
Notes on non-nominal systems:
Generator: run
Hours run: 12.5
From what time last night: 1758
To what time this morning: 0630
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 71 %
Diesel Reading – 39
Station Propane Reading – 57
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1462858
Water (static tank) (gal): 343
Static to Loft Pump used – yes
Water in Green Hab (gal): 267
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used:
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: no
Hours: 114.5
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 100
Currently charging:
Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Curiosity rover used: no
Hours: 119.6
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 100
Currently charging: yes
Notes on rovers: will have to juggle charging sequence.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General notes and comments: Weather day, no EVA ativity. HAB maintenance, mission planning for future EVAs and personal project/report time.
Summary of internet: No serious issues.
Summary of suits and radios: No issues.
Summary of Hab operations: Routine maintenance and housekeeping.
Summary of GreenHab operations: Routine watering.
Summary of Science Dome operations: Gypsum processing continued.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues remain unresolved.
Summary of any observatory issues: None, weathered out.
Summary of health and safety issues: HSO broke sim to address medical issue at Base Sopport.
Questions, concerns and requests to Mission Support: Need to replace 1 bottle of cough suppresent/expectorant consumed.